Write the quadratic equation whose roots are
3
and
−2
, and whose leading coefficient is
4
.
(Use the letter
x
to represent the variable.)

Answer by ikleyn(52778)

.
P(x) = 4*(x-3)*(x-(-2)) = 4*(x-3)*(x+2).